OpenStudy (anonymous):

The derivative of the x - axis is 0 , but what is the derivative of the y - axis ?

ganeshie8 (ganeshie8):

derivative = slope = undefined

OpenStudy (anonymous):

\[\frac{dy}{dx}=0 \\ \\ \frac{dx}{dy}=\frac{1}{0}\]